OpenStudy (anonymous):

The rate of intake, R, during respiratory cycle for a person at rest is proportional to a sine wave with period 6 seconds. If the rate of intake is 0.89 liters per second at 1.5 seconds, find the formula for R

OpenStudy (anonymous):

since time period is given as 6, we can calculate it's \[R(t)= 0.89 \sin( \frac{ \pi }{ 3 } * 1.5)\]

OpenStudy (anonymous):

@mnrydman

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

yep!!! it is.. you can solve further for values inside sin..but overall it's the equation. you know he equation will be of the form: r = c*sin(kt) right??

OpenStudy (anonymous):

yeah that makes sense

OpenStudy (anonymous):

awesome!! u just need to substitute the values ... since we know the period is six seconds, k must equal 2Pi/6 ...finally u will get the above R(t) @mnrydman

OpenStudy (anonymous):

\(\huge \cal \color{blue}{Clear?}\)